Country Able and Country Baker initially have the same real GDP per capita. Country Able experiences no economic growth, while Country Baker grows at a sustained rate of 7 percent. In 12 years, Country Baker's GDP will be approximately ___________ that of Country Able.

Question 14 options:

1)

triple

2)

double

3)

one-half

4)

one-fourth
Answer #1

Let initially their GDP was X

Each year growth of country baker=7% or 0.07

Given, able gdp is fixed at X.

After 12 year country baker GDP=X*(1+0.07)^12=X*(1.07)^12=X*2.25

Compare to baker gdp go able=2.25X/X=2.25

So baker has 2.25 times of able gdp after 12 years.

So baker has approximately double gdp that of able.

Option 2) is Right